OpenStudy (anonymous):

can any one help me with this problem? the position of a particle is given by the expression x=4.00cos(3.00pi(t)+pi) determine the frequency

OpenStudy (turingtest):

in the simple harmonic motion formula\[x(t)=x_0\cos(\omega t+\phi)\]omega is the angular velocity, phi is the phase shift, and x0 is the maximum displacement. uniform angular velocity is \[\omega=\frac{2\pi}T\]where T is the period of the motion. the frequency is given by\[f=\frac1T=\frac\omega{2\pi}\]